OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal...

OpenStudy (abbycross167):

A laborer can do a certain job in 5 days, a second can do the job in 6 days, and a third in 8 days. In what time can the three together do the job?

OpenStudy (abbycross167):

I did the steps once but my teacher said something was wrong...

OpenStudy (will.h):

explain your steps please

OpenStudy (abbycross167):

1/5+1/6+1/8=59/120 Combined rate * time (t) = 1 59 ------ t = 1 120 120 59 120 ------ * ------ t = ------ * 1 59 120 59 120 T= ------ = 2.0339 59 It will take them about two and a half days.

OpenStudy (abbycross167):

this is what my teacher said.... 120/59 does not equal 2 1/2 days!!! whoa!

OpenStudy (sshayer):

the teacher is correct 2 and half means 2.50etc

OpenStudy (mjdennis):

Sure, because 0.03 is way less than one-half. I think your math looks right, it is just your interpretation that is wrong..

OpenStudy (abbycross167):

So I should say it takes a little less than 2 1/2? @sshayer

OpenStudy (sshayer):

it is\[2\frac{ 2 }{ 59 }days\] or nearly 2 days or slightly more than 2 days

OpenStudy (sshayer):

2.03 is approx. 2 days

OpenStudy (abbycross167):

oh ok round down! I forgot! thank you!!!